OpenStudy (anonymous):

Prove that: if \[a + 2b + 3c = 7x\] then \[a^2 + b^2 + c^2 = (x-a)^2 + (2x-b)^2 + (3x-c)^2\]

OpenStudy (anonymous):

a+2b+3c=7x x=(a+2b+3c)/7 Put this into (x−a)^2+(2x−b)^2+(3x−c)^2 (x−a)^2+(2x−b)^2+(3x−c)^2 =(((a+2b+3c)/7)-a)^2 +(2(a+2b+3c)/7 -b)^2 +(3(a+2b+3c)/7 -c)^2 Then by expansion and grouping terms, we can get the answer.

OpenStudy (anonymous):

That seems way too long though.

OpenStudy (anonymous):

Here is how I did it